OpenStudy (anonymous):

verify: sec^2x/tanx+cot^2xtanx =tanx

OpenStudy (campbell_st):

well start by factoring the deominator \[\frac{\sec^2(x)}{\tan(x)(1 + \cos^2(x))} \]

OpenStudy (anonymous):

ok then you get the bottom to be tanx (cscx) right

OpenStudy (campbell_st):

oops should be \[\frac{\sec^2(x)}{\tan(x)(1 + \cot^2(x)}\] so yes you have \[\frac{1}{\tan(x)} \times \frac{\sec^2(x)}{\csc^2(x)} = \] so what do you think about sec^2/csc^2...?

OpenStudy (campbell_st):

\[\frac{\sec^2(x)}{\csc^2(x)} = \frac{1}{\cos^2(x)} \div \frac{1}{\sin^2(x)} = \frac{1}{\cos^2(x)} \times \frac{\sin^2(x)}{1} \] any thoughts..?

OpenStudy (campbell_st):

then multiply the answer above by \[\frac{1}{\tan(x)}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

is that right

OpenStudy (campbell_st):

thats correct you you could have just used \[\frac{1}{\tan(x)} \times \frac{\sin^2(x)}{\cos^2(x)} = \frac{1}{\tan(x)} \times \tan^2(x) = \tan(x)\]